OpenStudy (anonymous):

Suppose you draw a card from a deck of 52 cards. What is the probability that the card you draw is NOT an ace? There are 4 aces in the deck.

OpenStudy (diyadiya):

48/52

OpenStudy (stormfire1):

12/13

OpenStudy (anonymous):

52-4=48 48/52=12/13

OpenStudy (anonymous):

48/52=12/13 reduced

OpenStudy (anonymous):

1/13

OpenStudy (stormfire1):

@narutoboy: the question said what's the probability that it's NOT an ace.